# Hallucination Evaluation, Analysis and Mitigation in Multimodal Large Language Models

### v2

@@ 1 Introduction
POPE Li et al. (2023c) COCO Post-processed Acc. Co-occur.
HaELM Wang et al. (2023b) MS-COCO Post-processed Acc. Attention
MME Fu et al. (2024) Self-Sourced Manual Acc. -
AMBER Wang et al. (2024a) MS-COCO Post-processed Acc. -
MHaluBench Chen et al. (2024b) Self-Sourced Post-processed Prec. -
R-Bench Wu et al. (2024) COCO Post-processed Acc. Co-occur.
MMRel Nie et al. (2024) Visual-Genome Multi-source Acc. -
VALOR-EVAL Qiu et al. (2024) GQA Post-processed LLM-based Co-occur.
@@
Table 1: Comparsions of our proposed Reefknot benchmark with other relevant benchmarks.
In recent years, large language models (LLMs) have revolutionized the AI field by expanding their training data to trillions of tokens and increasing their parameter counts to hundreds of billions Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023). This has unlocked powerful emergent abilities, and seen widespread applications in diverse domains Achiam et al. (2023); Yan et al. (2024); Wang et al. (2023a). Recently, the community managed to combine visual backbones with powerful LLMs, resulting in multimodal large language models (MLLMs) Liu et al. (2023b). While this has led to advancements in multimodal scenarios, it also presents challenges for MLLMs, notably their tendency to generate hallucinations Zou et al. (2024). In LLMs, hallucinations occur when the model generates inaccurate or misleading factual information that cannot be supported by existing knowledge Zhang et al. (2023). However, such issues become more complex in MLLMs, as hallucinations can manifest as responses containing references or descriptions of the input image that are incorrect Bai et al. (2024); Huo et al. (2024a). Therefore, it is crucial to evaluate and mitigate these hallucinations to improve the trustworthy of MLLMs in real-world scenarios.
@@
To handle the aforementioned research gaps, we propose the first comprehensive benchmark Reefknot to evaluate the performance on relation hallucination. Unlike many benchmarks of MLLM that were constructed by automatic labeling technique, we construct the relation-based dataset based on semantic triplets retrieved from the scene graph dataset. Table 1 demonstrates that our triplets are from real-life scenarios, without any post-processing (e.g., segmentation and bounding box techniques), manual annotation, and synthetic method (e.g., diffusion-based generation). We categorize relation hallucinations into two types: perceptive, involving concrete relational terms like "on", "in", "behind" and cognitive, which encompasses more abstract terms such as "blowing" and "watching". Second, we evaluate the mainstream MLLMs on Reefknot via three diverse tasks across two types of relation hallucinations. Figure 2 also illustrates that relation hallucination can be more severe than object hallucination in current MLLMs, highlighting the importance of our evaluation. Furthermore, we propose a simple relation hallucination mitigation method named Detect-then-Calibrate. This originates from the experimental observation that when relation hallucinations occur, the response probability drops significantly, hovering just above 50% in extreme cases compared to the usual nearly 90%. Our method achieves an average improvement of 9.75% across three relation hallucination benchmarks. In summary, Our main contributions are as follows:
@@ 2 Relation Hallucination Benchmark
Table 3: Evaluation of hallucination rates on the different MLLMs. Additionally, we use bold to highlight the best performance of open-sourced MLLMs, and underline to emphasize the distinction between perception and cognition of LLaVA-13B.
@@ 3.2 Main results
We find the occurrence of cognitive hallucinations is generally lower than that of perceptive hallucinations, which may diverge from intuition. Across all models and settings, the incidence of perceptive hallucinations is consistently 10% higher than that of cognitive hallucinations. In the most extreme case, such as LLaVA-13B model in the MCQ setting, the rate of perceptive hallucinations is 30.16% higher than that of cognitive hallucinations. A possible reason for this phenomenon is that during the pre-training and fine-tuning processes, models often utilize large-scale image-caption datasets. These datasets typically contain detailed visual descriptions, enabling models to perform better in cognitive relationships such as running, eating, etc. Conversely, these models may struggle when dealing with some perceptive relationships based on common sense because they were ignored in the annotation process of the original dataset.
@@ 5 Detect-Then-Calibrate Mitigation Method
Note that $\lambda$ is the hyperparameter to control the degree of the intermediate layer. The algorithmic flow can be seen from the pseudo-code below:
@@ 6 Case study
To intuitively demonstrate the generation process of hallucination, we visualized a real case from Reefknot in Figure 8. In the image, when we ask “Is the boy eating pizza in the photo?", MLLMs are unable to provide precise answers. To investigate this, we analyze the probabilities associated with the top five predicted tokens at each layer of the language model. As illustrated in Figure 8, the model typically converges on the correct answer as it progresses from shallow to deep layers. Notably, hallucinations didn’t occur until the final layer. Additionally, we can observe that as the number of layers increases, the model increasingly focuses on narrowing down the range of answer choices, with the probability for both ‘yes’ and ‘no’ rising in deeper layers.
@@ 7 Related Work
Large Language Models (LLMs) are powerful tools, and exploring their interpretability Han et al. (2024); Huo et al. (2024b), operational mechanisms, and methods to ensure trustworthy responsesHu et al. (2023) has become a crucial area of research. In the realm of MLLMs, hallucinations are typically categorized into three distinct types: object, attribute, and relation, as outlined in prior studies Yin et al. (2024); Liu et al. (2024); Guan et al. (2024). At the object and attribute levels, a considerable number of representative benchmarks such as POPE Li et al. (2023c) and HaELM Wang et al. (2023b) have been introduced by researchers. Many benchmarks are accompanied by a diverse array of evaluation criteria, with generative criteria including CHAIR Li et al. (2023b), THRONE Kaul et al. (2024) and various discriminative criteria being particularly notable. The existing relation hallucinations benchmarks Wu et al. (2024); Nie et al. (2024) focus only on discriminative criteria. The gap of current benchmarks is to identify relationship pairs that accurately reflect the dynamic interactions.
Scene Graph Dataset
The concept of scene graphs was first introduced by Zhu et al. (2022), offering a novel method to describe and retrieve complex scenes within images. Building on this foundation, Krishna et al. (2016) further advanced the field by introducing the Visual Genome (VG) dataset in 2017. In recent years, there has been significant progress in leveraging this triplet-based structure for various applications. For instance, Yao et al. (2018) utilized scene graphs to generate more detailed image descriptions. Similarly, Li et al. (2019) enhanced VQA systems by incorporating scene graphs, improving the system’s ability to respond to complex questions about images.
@@
Confidence estimation Zou et al. (2023); Chen et al. (2024a) and calibration are essential for enhancing the reliability of LLMs such as GPT-3 Brown et al. (2020). To assess the confidence associated with outputs from LLMs, Kuhn et al. (2023) have developed a method called semantic entropy that utilizes linguistic invariances reflecting shared meanings. However, this method relies on accessing token-level probabilities, which are often unavailable through current black-box APIs. Kadavath et al. (2022) have designed prompts that encourage the models to self-assess their responses and to explicitly calculate the probability that an answer is true; while Lin et al. (2022) have prompted LLMs to provide both an answer and an accompanying confidence level. Furthermore, Manakul et al. (2023) proposed a sampling-based approach for detecting hallucinated facts in the outputs of models.
@@ A.2 Manual Checking Rules
A.3 More Error Analysis
@@ A.4 Prompt to Categorize
A.5 Potential concern
Why detect-than-calibrate method is specific to relation hallucination?
@@ A.5 Potential concern
What is the difference between detect-than-calibrate and contrastive decoding based methods like (i.e., vcd , dola)?
@@
A.6 VQA Criterion
@@
Table 7: Comparison of Hallucination Rates between DeBERTa-Based and GPT-Assisted Methods in VQA Settings
Here are the evaluation criteria for VQA questions. We use the DeBERTa [^1] https://huggingface.co/microsoft/deberta-v2-xlarge-mnli model to determine whether the models entail each other. Only when the label and response contain each other will it be judged as a correct reply. We show our key function code in Figure 15. To show potential bias associated with the form of constrained prompt and DeBERTa, we compared the evaluation results using a GPT-assisted method. In this setting, no restrictions were imposed on the prompt (e.g., "What is the relationship between A and B?"). The results, as shown in Table 7, demonstrate that when the responses are not constrained by the prompt, the performance of the vision language models in VQA about relationships is significantly impaired. To further validate the effectiveness of the DeBERTa model, we compared the judgments made by the DeBERTa with those of human checkers, which demonstrate that DeBERTa can accurately assess the correctness of the model’s responses in Table 8. Besides, we show a VQA case in Figure 17.
@@
sunlight is shining on train sunlight is illuminating train 1 1
bear is reading book bear is sitting on book 0 0
picture is hanging wall picture is on wall 1 1
man is lying on couch man is lying couch 1 1
Table 8: The comparison of results between DeBERTa and manual Check for real cases
A.7 More Cases
@@
A.8 Visualization of Relation Word
@@
3 #label is groundtruth;responese is the respone of VLM;

### v3

@@ 1 Introduction
POPE COCO Post-processed Acc. Co-occur.
HaELM MS-COCO Post-processed Acc. Attention
MME Self-Sourced Manual Acc. -
AMBER MS-COCO Post-processed Acc. -
MHaluBench Self-Sourced Post-processed Prec. -
R-Bench COCO Post-processed Acc. Co-occur.
MMRel Visual-Genome Multi-source Acc. -
VALOR-EVAL GQA Post-processed LLM-based Co-occur.
FAITHSCORE MS-COCO & LLaVA-1k Post-processed Faithscore Atomic Facts.
FIHA MS-COCO Foggy & VG Ruled-based Acc. Annotation Free
TIFA Synthetic Caption-based Acc. T2I Evaluation
@@
Table 1: Comparsions of our proposed Reefknot benchmark with relevant benchmarks. POPE Li et al. (2023c), HAELM Wang et al. (2023b), MME Fu et al. (2024), AMBER Wang et al. (2024a), MHaluBench Chen et al. (2024b), R-bench Wu et al. (2024), MMRel Nie et al. (2024), VALOR-EVAL Qiu et al. (2024), FAITHSCORE Jing et al. (2024), FIHA Yan et al. (2024a), TIFA Hu et al. (2023b)
In recent years, large language models (LLMs) have revolutionized the AI field by expanding their training data to trillions of tokens and increasing their parameter counts to hundreds of billions Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023). This has unlocked powerful emergent abilities, and seen widespread applications in diverse domains Achiam et al. (2023); Yan et al. (2024b); Wang et al. (2023a). Recently, the community managed to combine visual backbones with powerful LLMs, resulting in multimodal large language models (MLLMs) Liu et al. (2023b). While this has led to advancements in multimodal scenarios, it also presents challenges for MLLMs, notably their tendency to generate hallucinations Zou et al. (2024). In LLMs, hallucinations occur when the model generates inaccurate or misleading factual information that cannot be supported by existing knowledge Zhang et al. (2023). However, such issues become more complex in MLLMs, as hallucinations can manifest as responses containing references or descriptions of the input image that are incorrect Bai et al. (2024); Huo et al. (2024a). Therefore, it is crucial to evaluate and mitigate these hallucinations to improve the trustworthy of MLLMs in real-world scenarios.
@@
To handle the aforementioned research gaps, we present a specific definition for relation hallucination and propose the first comprehensive benchmark Reefknot to evaluate the performance on relation hallucination. Relation hallucination refers to the phenomenon in which MLLMs misinterpret the logical relationships between two or more objects in the corresponding image. Unlike many benchmarks of MLLM that were constructed by automatic labeling technique, we construct the relation-based dataset based on semantic triplets retrieved from the scene graph dataset. Table 1 demonstrates that our triplets are from real-life scenarios, without any post-processing (e.g., segmentation and bounding box techniques), manual annotation, and synthetic method (e.g., diffusion-based generation). We categorize relation hallucinations into two types: perceptive, involving concrete relational terms like "on", "in", "behind" and cognitive, which encompasses more abstract terms such as "blowing" and "watching". Second, we evaluate the mainstream MLLMs on Reefknot via three diverse tasks across two types of relation hallucinations. Figure 2 also illustrates that relation hallucination can be more severe than object hallucination in current MLLMs, highlighting the importance of our evaluation. Furthermore, we propose a simple relation hallucination mitigation method named Detect-then-Calibrate. This originates from the experimental observation that when relation hallucinations occur, the response probability drops significantly, hovering just above 50% in extreme cases compared to the usual nearly 90%. Our method achieves an average improvement of 9.75% across three relation hallucination benchmarks. In summary, Our main contributions are as follows:
@@ 2 Relation Hallucination Benchmark
Table 3: Evaluation of hallucination rates on the different MLLMs. Additionally, we use bold to highlight the best performance of open-sourced MLLMs, and underline to emphasize the distinction between perception and cognition of LLaVA-13B. Note all experiments are done with the temperature to 0 to keep the reproducibility.
@@ 3.2 Main results
We find the occurrence of cognitive hallucinations is generally lower than that of perceptive hallucinations, which may diverge from intuition. Across all models and settings, the incidence of perceptive hallucinations is consistently 10% higher than that of cognitive hallucinations. In the most extreme case, such as LLaVA-13B model in the MCQ setting, the rate of perceptive hallucinations is 30.16% higher than that of cognitive hallucinations. This phenomenon may be caused by the fact that models often utilize large-scale image-caption datasets during the pre-training and fine-tuning processes. These datasets typically contain detailed visual descriptions, enabling models to perform better in cognitive relationships such as running, eating, etc. Conversely, these models may struggle when dealing with some perceptive relationships based on common sense because they were ignored in the annotation process of the original dataset.
@@ 5 Detect-Then-Calibrate Mitigation Method
Note that $\lambda$ is the hyperparameter to control the degree of the intermediate layer; $\mathbf{r}$ is the token generated after calibration; $\alpha$ represents the degree of calibration operation. The algorithmic flow can be seen from the pseudo-code below:
@@ 6 Case study
To intuitively demonstrate the generation process of hallucination, we visualized a real case from Reefknot in Figure 8 Nostalgebraist (2020). In the image, when we ask “Is the boy eating pizza in the photo?", MLLMs are unable to provide precise answers. To investigate this, we analyze the probabilities associated with the top five predicted tokens at each layer of the language model. As illustrated in Figure 8, the model typically converges on the correct answer as it progresses from shallow to deep layers. Notably, hallucinations didn’t occur until the final layer. Additionally, we can observe that as the number of layers increases, the model increasingly focuses on narrowing down the range of answer choices, with the probability for both ‘yes’ and ‘no’ rising in deeper layers.
@@ 7 Related Work
Large Language Models (LLMs) are powerful tools, and exploring their interpretability Han et al. (2024); Huo et al. (2024b), operational mechanisms, and methods to ensure trustworthy responsesHu et al. (2023a) has become a crucial area of research. In the realm of MLLMs, hallucinations are typically categorized into three distinct types: object, attribute, and relation, as outlined in prior studies Yin et al. (2024); Liu et al. (2024); Guan et al. (2024). At the object and attribute levels, a considerable number of representative benchmarks such as POPE Li et al. (2023c) and HaELM Wang et al. (2023b) have been introduced by researchers. Many benchmarks are accompanied by a diverse array of evaluation criteria, with generative criteria including CHAIR Li et al. (2023b), THRONE Kaul et al. (2024) and various discriminative criteria being particularly notable. The existing relation hallucinations benchmarks Wu et al. (2024); Nie et al. (2024) focus only on discriminative criteria. The gap of current benchmarks is to identify relationship pairs that accurately reflect the dynamic interactions.
@@
Confidence estimation Zou et al. (2023); Chen et al. (2024a) and calibration are essential for enhancing the reliability of LLMs such as GPT-3 Brown et al. (2020). To assess the confidence associated with outputs from LLMs, Kuhn et al. (2023) have developed a method called semantic entropy that utilizes linguistic invariances reflecting shared meanings. However, this method relies on accessing token-level probabilities, which are often unavailable through current black-box APIs. Kadavath et al. (2022) have designed prompts that encourage the models to self-assess their responses and to explicitly calculate the probability that an answer is true; while Lin et al. (2022) have prompted LLMs to provide both an answer and an accompanying confidence level.
@@ 9 Limitation
Acknowledgement
This work was supported by Open Project Program of Guangxi Key Laboratory of Digital Infrastructure (Grant Number: GXDIOP2024015); Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
@@ A.2 Manual Checking Rules
A.3 Complete guideline for Manual Check
Next I will provide you with three types of questions and corresponding images. The types of questions are Multiple Choice Question (MCQ), Yes/No(Y/N), and Visual Question Answer (VQA). MCQ asks about the direct relationship between two objects that already exist in the picture, and the question presupposes four options, with only one option being the final correct answer; Y/N questions presuppose a relationship between the two objects in the image, with the objects already existing in the image; and the VQA is the same as the multiple-choice question setup, but with no options provided. Your task is to make judgments about the reasonableness of the question setup and the relevance of the question to the picture.
The degree of reasonableness of the question refers to the fact that the objects involved in the question set must be present in the picture and that the collocation between the two objects has a certain degree of immobility; typical fixed-type collocations are triplets like, <ear on face>, <hair on head>, and when you come across these kinds of relational words in the question set, you should filter them out straight away.
The relevance of the picture means that the objects appearing in the question should occupy the main part of the picture, if the picture appears in the minutiae of the question (e.g., when it comes to some very subtle information about the background), you should exclude the picture and the question.
You can modify, discard and keep the pairs you see, with the following precautions:
  1. Please exclude unusual and meaningless semantic triplets like <man, wearing, shirt>.
  2. Please carefully check the question setting and rewrite some question setting to keep diversity.
  3. Please make a final judgment on each pair, you can choose to keep, change and keep, or delete.
A.4 More Error Analysis
@@ A.5 Prompt to Categorize
A.6 Potential Concerns
❶ Why detect-than-calibrate method is specific to relation hallucination?
@@ A.6 Potential Concerns
❷ What is the difference between detect-than-calibrate and contrastive decoding based methods like (i.e., vcd , dola)?
@@
❸ Why no more experiments on other benchmarks (i.e., POPE , CHAIR)?
POPE Li et al. (2023c) and CHAIR Li et al. (2023b) are indeed well-known benchmarks for hallucination evaluation. However, POPE is specifically focused on object-level hallucinations, while CHAIR is a generative metric for evaluating object-level hallucinations (caption-based).
Our Detect-Than-Calibrate focuses on exploring the mechanisms and mitigation of relation-level hallucinations. In Figure 6, we demonstrate the hallucination ratio of POPE under different entropy distributions. As shown in Figure 6, POPE does not show an increase in hallucinations with higher entropy, making it unsuitable for evaluating our entropy-based Detect-than-Calibrate method. CHAIR, on the other hand, is a generative metric for object-level hallucinations, whereas our method targets discriminative relation-level hallucination mitigation. Therefore, it is not feasible to evaluate our approach using POPE and CHAIR.
To validate the effectiveness of our method, besides experiments on our Reefknot, we also conducted experiments on two benchmarks specifically targeting relation-level hallucination, R-bench Wu et al. (2024) and MMrel Nie et al. (2024), in addition to our own Reefknot. The results in Table 5 demonstrate the effectiveness of our approach. Additionally, in the paper, we explore the distinction between object-level and relation-level hallucinations in Figures 26, and Table 6.
❹ The Proof Reliability of DebERTa
following the setting from Kuhn et al. (2023); Farquhar et al. (2024), we reformulated VQA task as an NLI task by appending the instruction “Please answer in the following format: Subject is <relation> Object" to the prompt. This raises two potential concerns: First, there is a concern about DeBERTa’s ability to distinguish between MLLM’s response and the ground truth; Second, there is a concern about how the introduction of answer templates might influence the results.
To address the concern regarding DeBERTa’s capabilities, we selected a number of examples from the study for verification and had human annotators annotate these examples for comparison, as shown in Table 8. The results showed that over 95% of the cases yielded identical annotations between DeBERTa and the human annotators, demonstrating that DeBERTa is well-suited for this task.
To address the concern about the introduction of answer templates, we compared responses with and without templates. Responses generated with the template were structured as phrases, while those without the template were more open-ended and variable. We employed GPT-4o-mini to assess whether the open-ended responses were semantically equivalent to the ground truth. Table 7 shows that without a constrained template, MLLMs tend to show a higher hallucination rate, which proves the necessity of the template in prompts.
❺ Why do you set $E(X)>$ 0.6 as a threshold to calibrate? The ratio seems unchanged between $E(X)$=0.4 and $E(X)$=0.6 in Figure 6.
The higher the entropy value, the lower the confidence of the response, and the greater the likelihood of relation hallucinations occurring. The $E(X)$ values we use serve as a threshold for our calibration operation. Calibration is performed only when the entropy exceeds this threshold. The benefit of this approach is that it helps avoid unnecessary calibration for cases that do not involve hallucinations. We chose $E=0.6$ as a threshold to focus on cases where hallucinations are more likely to occur, ensuring that calibration is applied only to cases with a higher degree of certainty of hallucination, and try our best to reduce calibration on non-hallucination cases.
❻ What is the meaning of meaningless questions in manual verification when constructing Reefknot?
"Meaningless questions" are defined as questions like “Are there ears on a milk cow?” or “Are there eyes on a man?” These types of triplets are fixed combinations, and even without visual information, they can easily be inferred through common sense.
A.7 VQA Criterion
@@
Table 7: Comparison of Hallucination Rates between DeBERTa-Based and GPT-Assisted Methods in VQA settings.
Here are the evaluation criteria for VQA questions. We use the DeBERTa [^1] https://huggingface.co/microsoft/deberta-v2-xlarge-mnli model to determine whether the models entail each other. Only when the label and response contain each other will it be judged as a correct reply. We show our key function code in Figure 15. To show potential bias associated with the form of constrained prompt and DeBERTa, we compared the evaluation results using a GPT-assisted method. In this setting, no restrictions were imposed on the prompt (e.g., "What is the relationship between A and B?"). The results, as shown in Table 7, demonstrate that when the prompt does not constrain the responses, the performance of MLLMs in VQA about relationships is significantly impaired. To further validate the effectiveness of the DeBERTa model, we compared the judgments made by the DeBERTa with those of human checkers, which demonstrate that DeBERTa can accurately assess the correctness of the model’s responses in Table 8. Besides, we show a VQA case in Figure 17.
@@
sunlight is shining on train sunlight is illuminating train
bear is reading book bear is sitting on book
picture is hanging wall picture is on wall
man is lying on couch man is lying couch
dog is barking at stranger dog is at stranger
rain is falling on roof rain is on roof
car is parked in garage car is in garage
fish is swimming pond fish is darting water
dog is barking yard dog is howling garden
fire is burning hearth fire is crackling fireplace
Table 8: The comparison of results between DeBERTa and Manual Check for real cases from our Reefknot.
A.8 More Cases
@@
A.9 Visualization of Relation Word
@@
3 #label is groundtruth; response is the reply of VLM;